hi can you help me solve this twp step equation: n/3 +6= 22
Answer by mananth(16946)

n/3 +6 =22
add -6
n/3 +6 -6 = 22-6
n/3 = 16
multiply by 3
n/3 * 3 = 16*3
n =48
...
m.ananth@hotmail.ca